OpenStudy (anonymous):

Evaluate the expression without the aid of a calculator. arctan 0 arccos 1

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

HELP PLEASE!!! :(

OpenStudy (anonymous):

tg x = 0 find x: tg x = sinx/cos x sinx =0 for x = nPi

OpenStudy (anonymous):

just like in inverse functions with logarithms and exponentials, you rewrite what this means in it's inverse form: \[\huge arctan0=y\Leftrightarrow tany=0 \] what is y?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

y=x

OpenStudy (anonymous):

is this your question? \[\huge arctan0 \cdot arcos1 \]????

OpenStudy (anonymous):

THEY JUST WANT YOU TO EVALUATE IT

OpenStudy (anonymous):

ow... my ears!

OpenStudy (anonymous):

virtual ears...

OpenStudy (anonymous):

lol sorry my caps key is having issues

OpenStudy (anonymous):

ok... so arctan0 = 0 is correct...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

arctg 0 = n*Pie arcos 1 = 2n*Pie

OpenStudy (anonymous):

you can do that with arccos1=y ... you'll need to rewrite it in it's inverse form: cosy = 1 ...what's y?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

arctan is multivalued function....

OpenStudy (anonymous):

0 is just one solution...

OpenStudy (anonymous):

yes, irealize that @myko, but we're just looking for the principal value for these inverse trig functions..

OpenStudy (anonymous):

and the range for arctan is restricted to -pi/2 to pi/2

OpenStudy (anonymous):

ok i guess the point is Arctan and arctan.

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

i think he just wants the value for arctan0, and not the solutions to the equation 0=arctanx

OpenStudy (anonymous):

the equation will have the solutions you pointed out... but yeah, some texts interchange Arctan an arctan without giving a second thought... good point @myko ....:)

OpenStudy (anonymous):

so arccos1=y means cosy = 1.... what was y again?